OpenStudy (anonymous):

Write the equation of the line perpendicular to 9x+2y=18 and passes through the point (1,4). Possible answers- a) 9x-2y=1, b) 2x-9y=-34, c) 2x+9y=38, d) 9x+2y=17

OpenStudy (nikato):

do you know how to find the slope of 9x + 2y=18

OpenStudy (anonymous):

no

OpenStudy (nikato):

do you recognize the point-slope form? y=mx+b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

yes so would i not put anything in the slope spot?

OpenStudy (nikato):

no. you have to do a little algebra such that it would look like that form

OpenStudy (nikato):

so lets start by subtracting 9x on both sides. what do you have now?

OpenStudy (anonymous):

I'm not sure what the set up of the equation is

OpenStudy (nikato):

9x +2y=18 -9x -9x -------------- 2y=-9x +18 does this make sense to you?

OpenStudy (anonymous):

yes

OpenStudy (nikato):

now divide both sides by 2 so that y is isolated on the left side

OpenStudy (anonymous):

Wouldn't that cancel out the 2?

OpenStudy (nikato):

correct 2y=-9x +18 -- ------- 2 2 y=-9/2x +9 do you see why i did this? and doesnt this look familiar now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I don't recall

OpenStudy (nikato):

i meant slope-intercept form y=mx+b

OpenStudy (anonymous):

oh yes

OpenStudy (nikato):

so whats the slope?

OpenStudy (anonymous):

9?

OpenStudy (nikato):

\[y=-\frac{ 9 }{ 2 }x +9\]

OpenStudy (nikato):

what's the slope?

OpenStudy (anonymous):

I don't know. I don't understand this

OpenStudy (nikato):

whats the "m"

OpenStudy (anonymous):

I just said I don't know

OpenStudy (nikato):

its -9/2

OpenStudy (nikato):